If combined age of family is 81 and the father is 3 years older than mother and mother is 27 years older from son then find the age of all three separatly

Answer by lwsshak3(11628)

If combined age of family is 81 and the father is 3 years older than mother and mother is 27 years older from son then find the age of all three separatly
let x= age of son
x+27=age of mother
x+27+3= age of father
x+x+27+x+27+3=81
3x=24
x=8
age of son=8
age of mother=35
x+27+3= age of father=38